OpenStudy (anonymous):

In september 1998 the population of the country of West Goma in millions was modeled by f(x) = 16.7e^0.0006x. At the same time the population of East Goma in millions was modeled by g(x) = 13.7e^0.0187x. In both formulas x is the year, where x=0 corresponds to September 1998. Assuming these trends continue, estimate the year when the population of West Goma will equal the population of East Goma.

OpenStudy (anonymous):

you have said to estimate x when both populations are equal... so, it is assumed that both populations are equal and written as 16.7e^0.0006x=13.7e^0.0187x The entire equation is divided by 13.7e^0.0187x which gives, 16.7e^0.0006x/13.7e^0.0187x=1 splitting the terms (16.7/13.7)(e^0.0006x/e^0.0187x)=1 16.7/13.7=1.21 e^0.0006x/e^0.0187x=e^-0.0181x so we get, 1.21e^-0.0181x=1 divide the whole equation by 1.21, e^-0.0181x=0.8264 taking natural logarithm on both sides, -0.0181x= -0.190676361 => x=10.53 x cannot be decimal number so approximating it to the closest integer 11 so aftr 11 yrs from 1998 population will be equal that is at 2009 population will be equal